# Kudla Millson lift of toric cycles and restriction of Hilbert modular forms

### v3

@@ 5.  Spans of diagonal restrictions and toric cycles
Let $(V,Q)$ be a rational quadratic space of even dimension and signature $(p,q)$ with $p\geq q>0$. Let $\varphi=\mathbf{1}_{\widehat{L}}$ be the characteristic function of a lattice $\widehat{L}\subset V_{\widehat{\mathbb{Q}}}$, such that $L=\widehat{L}\cap V$ is even and of level $1$ (unimodular). By Poincaré duality, the pairing
@@
is non-degenerate. For a subspace $V\subset H^{q}(Y,\mathbb{C})$ let $V^{\perp}\subset H_{q}(Y,\mathbb{C})$ denote the orthogonal complement with respect to the pairing. If we suppose that $q$ is odd, then the Kudla-Millson lift is in fact a lift
$\Theta=\Theta_{\varphi}\colon H_{q}(Y,\mathbb{C})\longrightarrow S_{d}(\operatorname{SL}_{2}(\mathbb{Z}))$ (5.2)
into the space of cusp forms of weight $d=\frac{p+q}{2}$; see [KM90, Theorem. 2]. On the other hand, the Kudla-Millson lift has an adjoint
@@
We have $p+q>4=\max(4,3+r)$ where $r$ is the Witt index. Moreover, we assumed that $L$ is an even unimodular lattice. Hence, the injectivity of $\overline{\Theta}$ is the content of Corollary 1.2 of [BF10]. Suppose that $f\in S_{d}(\operatorname{SL}_{2}(\mathbb{Z}))$ is a nonzero cusp form that is not contained in the image of $\Theta$. Without loss of generality we can assume that $f$ is orthogonal to $\operatorname{Im}(\Theta)$ with respect to the Peterson inner product, otherwise we replace $f$ by $f-\textrm{proj}(f)$ where $\textrm{proj}(f)$ is the orthogonal projection of $f$ onto $\operatorname{Im}(\Theta)$ with respect to the Peterson inner product. Hence, for every $C\in H_{q}(Y,\mathbb{C})$ we have
@@
Let $S$ be the set of characters $\chi\colon\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})\longrightarrow\mathbb{C}^{\times}$ with the same assumptions as in the rest of the paper. We define twe following two subspaces. First let
@@
The result follows by taking the difference of (5) and (5.13) and using that $\ker(\Theta)=\operatorname{Im}(\overline{\Theta})^{\perp}=H_{\operatorname{cycle}}^{\perp}$. The last equality is due to Kudla and Millson [KM88, Theorem. 4.2]. We recall the proof. First, if $C\in H_{\operatorname{cycle}}^{\perp}$, then $C\in\ker(\Theta)$ since the Fourier coefficients of $\Theta(C)$ are $\langle C,C_{n}(\varphi)\rangle=0$. Hence, we have $H_{\operatorname{cycle}}^{\perp}\subset\ker(\Theta)$. On the other hand, consider the $n$-th Poincaré series of weight $d$ defined

### v4

@@ 5. Spans of diagonal restrictions and toric cycles
Let $(V,Q)$ be a rational quadratic space of even dimension and signature $(p,q)$ with $p\geq q>0$. Let $\varphi=\mathbf{1}_{\widehat{L}}$ be the characteristic function of a lattice $\widehat{L}\subset V_{\mathbb{A}_{f}}$, such that $L=\widehat{L}\cap V$ is even and unimodular. By Poincaré duality, the pairing
@@
is non-degenerate. For a subspace ${U}\subset H^{q}(Y,\mathbb{C})$ let ${U}^{\perp}\subset H_{q}(Y,\mathbb{C})$ denote the orthogonal complement with respect to the pairing. If we suppose that $q$ is odd, then the Kudla-Millson lift is in fact a lift
$\Theta=\Theta^{V}_{\varphi}\colon H_{q}(Y,\mathbb{C})\longrightarrow S_{d}(\operatorname{SL}_{2}(\mathbb{Z}))$ (5.2)
into the space of cusp forms of weight $d=\frac{p+q}{2}$; see [17, Theorem. 2]. On the other hand, the Kudla-Millson lift has an adjoint
@@
We have $p+q>4=\max(4,3+r)$ where $r$ is the Witt index. Moreover, we assumed that $L$ is an even unimodular lattice. Hence, the injectivity of $\overline{\Theta}$ is the content of Corollary 1.2 of [6]. Suppose that $f\in S_{d}(\operatorname{SL}_{2}(\mathbb{Z}))$ is a nonzero cusp form that is not contained in the image of $\Theta$. Without loss of generality we can assume that $f$ is orthogonal to $\operatorname{Im}(\Theta)$ with respect to the Peterson inner product, otherwise we replace $f$ by $f-\textrm{proj}(f)$ where $\textrm{proj}(f)$ is the orthogonal projection of $f$ onto $\operatorname{Im}(\Theta)$ with respect to the Peterson inner product. Hence, for every $C\in H_{q}(Y,\mathbb{C})$ we have
@@
[…]
@@
The result follows by taking the difference of (5) and (5.13) and using that $\ker(\Theta)=\operatorname{Im}(\overline{\Theta})^{\perp}=H_{\operatorname{cycle}}^{\perp}$. The last equality is due to Kudla and Millson [16, Theorem. 4.2]. We recall the proof. First, if $C\in H_{\operatorname{cycle}}^{\perp}$, then $C\in\ker(\Theta)$ since the Fourier coefficients of $\Theta(C)$ are $\langle C,C_{n}(\varphi)\rangle=0$. Hence, we have $H_{\operatorname{cycle}}^{\perp}\subset\ker(\Theta)$. On the other hand, consider the $n$-th Poincaré series of weight $d$ defined